OpenStudy (anonymous):

Rationalize the denominator of 5/(9 + √5)

OpenStudy (anonymous):

\[\frac{5}{9+\sqrt{5}}\times\frac{9-\sqrt{5}}{9-\sqrt{5}}=\frac{45-5\sqrt{5}}{81-9\sqrt{5}+9\sqrt{5}-5}=\frac{45-5\sqrt{5}}{76}\]

OpenStudy (anonymous):

put 6x3/4 in radical form